# Quantum optical neural networks with programmable nonlinearities

### v1

@@ II.2 Bell-state discrimination
$|\Psi^{(\pm)}\rangle=\frac{|0\rangle_{L1}|0\rangle_{L2}+|1\rangle_{L1}|1\rangle_{L2}}{\sqrt{2}}=\frac{|0101\rangle\pm|1010\rangle}{\sqrt{2}},$
$|\Phi^{(\pm)}\rangle=\frac{|0\rangle_{L1}|1\rangle_{L2}+|1\rangle_{L1}|0\rangle_{L2}}{\sqrt{2}}=\frac{|0110\rangle\pm|1001\rangle}{\sqrt{2}},$ (7)
where the right-hand side are the Bell-states encoded by $N=2$ photons and $M=4$ modes. The discriminator can transform each of the Bell-states $|\Psi^{(in)}_{j}\rangle\in\{|\Psi^{(+)}\rangle,|\Psi^{(-)}\rangle,|\Phi^{(+)}\rangle,|\Phi^{(-)}\rangle\}$ into a multimode Fock state $|\Psi^{(out)}_{j}\rangle\in\{|0101\rangle,|1010\rangle,|0110\rangle,|1001\rangle\}$. In turn, measuring these Fock states by photodetection unambiguously tells about a corresponding input state.
@@
$|\Theta^{(\pm)}\rangle=\frac{|1100\rangle\pm|0011\rangle}{\sqrt{2}},$ (8)
which are equally useful. Although states (8) do not respect the dual-rail encoding, they can still serve as a maximally entangled resource that fueled a quantum algorithm, provided cheap linear transformation is performed. Namely, to convert these states into any of the Bell states, one can apply simple two-mode operations, implemented deterministically by linear optics [34]. Therefore, in the following we consider the states (8) on par with the Bell states (II.2). We demand that the discriminator convert the input states $|\Theta^{(+)}\rangle$ and $|\Theta^{(-)}\rangle$ into $|\Psi^{(out)}_{5}\rangle=|1100\rangle$, $|\Psi^{(out)}_{6}\rangle=|0011\rangle$, respectively.
@@
Figure 8: Bell-state discrimination performed by the QONNs: state infidelity $1-F_{disc}$ as a function of circuit depth $D$ for four- (a) and six-state discrimination (b) at different static phase-shift biases $\varphi_{b}=0$ of the QONN programmed by nonlinearities (solid curves). The dependencies corresponding to the QONN programmed by linear optics is plotted by the dashed curves.
Figure 9: Distribution of the parameters of the trained QONNs: values of the NMZI nonlinearities and the MZI phase-shifts of the four-state (a) and six-state (b) discriminator at $\varphi_{b}=0$ with single-qubit corrections. The parameters correspond to the QONNs with $D=1$ (a) and $D=5$ (b) (see Fig. 8).
@@
The results suggest that the NMZI meshes, both with and without single-qubit corrections, can perform state discrimination provided proper phase-shift biases are set. The lowest depths at which the QONNs implement perfect discriminators are achieved at $\varphi_{b}=0$, corresponding to the identity operation of the NMZI in the linear regime. Specifically, depths of $D=1$ and $D=5$ are sufficient for the four- and six-state discrimination, respectively (with single-qubit corrections). A phase-shift of $\varphi_{b}=\pi$ (corresponding to the mode-swapping operation in the linear regime of NMZI) also works but requires much larger depths: $D=3$ and $D=11$ for the four- and six-state discrimination, respectively.
Training only the variational core without the input and output MZIs works as well, however, this comes at increasing the depths even further. In Fig. 8 we plotted the infidelities without the single-qubit corrections for $\varphi_{b}=0$ and $\varphi_{b}=\pi$. For these biases, it is possible to achieve convergence of the optimization algorithm only for $\varphi_{b}=\pi$, because NMZI mode swapping operation allows to redistribute the photons between modes. The minimal circuit depths in this case are $D=19$ and $D=23$.
@@
We summarize the relevant parameters of the QONN performing the state discrimination task in Table 3. Notably, our QONN with adjustable nonlinearities is less demanding than QONNs with static nonlinearities, as evident from Fig. 8. Specifically, the QONN trained with linear optics requires $D=1$ and $D=3$ layers to achieve high-quality discrimination of the four- and six-state systems, respectively. Our QONN requires fewer adjustable nonlinearities than its counterpart trained with linear optics to find the optimal circuits. Additionally, our approach could find the QONN can operate as a four-state discriminator at a lower depth of $D=1$.
Table 3: Circuit depth $D$ and parameter count $P$ of the QONNs peforming state discrimination (from the results shown in Fig. 8).
@@
cell11 = r=2, vlines, hline1,3,7 = -, hline2 = 2-5, & $D$ $P$
$4$-/$6$-state $4$-/$6$-state
NL ($\varphi_{b}=0$) 1/5 2/14
NL ($\varphi_{b}=\pi$) 3/11 8/32
NL ($\varphi_{b}=\pi$, no MZI) 19/23 56/68
LO 1/3 24/48
@@ II.3 Variational quantum algorithms
We next consider QONN performing a variational quantum algorithm (VQA). VQAs enable solving certain types of computational problems by mapping them into corresponding Hamiltonians $\hat{H}$, whose ground state encodes the solutions. In order to find a ground state by VQA, one prepares parametrized trial states $|\psi(\bm{\chi},\bm{\theta})\rangle$ on a quantum computer and estimates the Hamiltonian expectation value by measuring $E(\bm{\chi},\bm{\theta})=\langle\psi(\bm{\chi},\bm{\theta})|\hat{H}|\psi(\bm{\chi},\bm{\theta})\rangle$ — the cost function to be optimized by the classical computer, which global minimum $E_{\text{min}}\leq{}E(\bm{\chi},\bm{\theta})$ correspond to a solution. In contrast to implementing specific quantum operations, such as those described above, solving problems with the variational quantum eigensolver (VQE) does not necessitate the exact generation of a specified state. VQE may provide good approximate solutions for a class of prepared quantum states not necessarily coinciding with the true ground state. Despite this relaxed complexity, solving practical problems by VQA is coupled with preparation of large-qubit quantum states, which, on the one hand, is too sophisticated due to the limitations of the current NISQ devices and, on the other hand, hard to optimize because of the unsolved issue of barrens plateau [29]. Therefore, developing less demanding quantum architectures for VQE is required.
@@
where $\hat{X}_{i},\hat{Y}_{i},\hat{Z}_{i}$ are the Pauli operators describing spin with index $i$. A particular Hamiltonian is specified by the interactions between the spins described by interaction coefficients $J_{ij}^{(\mu)}$ ($\mu=x,y,z$) with interaction graphs $G^{(\mu)}$ and external fields $h_{i}$. Given specified interactions and external fields, VQA boils down to finding the Hamiltonian ground state energy using an optimization algorithm run on a classical computer. Depending on the interaction graphs of the Hamiltonians, various quantum systems with different properties can be catched [39, yazyev2008magnetic, 40]
@@
Figure 11: VQE of the Heisenberg Hamiltonian performed by the QONNs: energy error $\Delta{}E=E_{\text{exact}}-E_{\text{qonn}}$ as a function of depth $D$ of the QONN programmed by nonlinearities at different values the NMZI phase-shift bias $\varphi_{b}$ (solid curves) and the dependence for the QONN programmed by linear optics (dashed curve). Each dependence correspond to $15$ Hamiltonians, obtained by randomly generating $J_{1}$, $J_{2}$ and $h_{i}$ from a uniform distribution in the range from 00 to $1$.
Specifically, we sampled a number of Hamiltonians by generating values of $J_{1}$, $J_{2}$ and $h_{i}$ ($i=\overline{1,5}$) randomly from the uniform distribution over in the range $[0,1]$. Here, we quantify the infidelity by the energy error $\Delta{}E=E_{\text{exact}}-E_{\text{qonn}}$, where $E_{\text{qonn}}$ is the energy obtained by training/optimizing a QONN, $E_{\text{exact}}$ is its exact value obtained by the diagonalization of the Hamiltonian matrix using a standard library (Python SciPy). The obtained dependencies of error on the circuit depth $D$ is shown in Fig. 11. As before, not all static phase-shift values $\varphi_{b}$ enabled find as a high-accuracy solution. As we see, the value $\varphi_{b}=0$ (corresponding to the mode swapping operation in the linear regime of NMZI) does not yield satisfactory quality. At the same time, all the rest of the values work almost equally good. According to Fig. 11, it requires minimum $D=7$ layers for the best case of $\varphi_{b}=\pi/2$ for the QONN to be able to find high-quality solutions. The QONN programmed by linear optics, for which the simulation results are also plotted in Fig. 11, more beneficial in terms of the minimal circuit depth, as $D=1$ is enough to achieve a high-quality solution. However, counting the number of programmable parameters suggests that our QONN is advantageous. Indeed, our QONN required $P_{\text{NL}}=62$ programmable nonlinearities, while the counterpart QONN necessitates using almost three times more programmable phase-shifts $P_{\text{LO}}=180$.
Table 4: Circuit depth $D$ and parameter count $P$ of the QONNs peforming VQE of the Heisenberg Hamiltonian (see Fig. 11). NL and LO are the QONNs with adjustable nonlinearities and the counterparts trained by linear optics, respectively.
@@
vlines, hline1-2 = -, hline4 = -, & $D$ $P$
NL ($\varphi_{b}=\pi/2$) 7 62
LO 1 180
It is worth noting that the circuit depth, and therefore the number of adjustable parameters in QONNs, can be further reduced by using different building blocks. Specifically, the ability of a QONN layer to implement non-local interactions may be more advantageous for certain problems compared to the two-mode interactions of NMZI layers [43]. Our proposal is not confined to this specific QONN architecture. We suggest exploring the use of programmable nonlinearities in other architectures that may be developed in the future.
@@ II.4 Training complexity
Figure 12: Training the QONNs for preparation of the $4$-qubit GHZ-state: a) the set of infidelities to which the optimization algorithm has converged in $30$ optimization runs and the corresponding number of evaluations of the cost function; b) examples of convergence curves in $10$ optimization runs. The parameters of the QONNs correspond to Fig. 5e) and Fig. 5f) at $D=9$ and $D=3$, respectively.
@@ III Discussion and Outlook
We propose a novel approach for constructing parametrized quantum optical circuits using programmable nonlinearities. Using NMZI meshes as an example of such programmable circuits, we demonstrate that high-quality operation can be achieved by training these nonlinearities. The key insight of our work is that relying solely on programmable linear optical interference is suboptimal for implementing QONNs. Our approach, which incorporates programmable nonlinearities, significantly reduces the number of adjustable parameters and even may decrease the circuit depth, thereby decreasing errors from circuit components and resulting in more compact designs. Additionally, variational circuits can benefit from a lower number of adjustable parameters, allowing for more computationally efficient optimization algorithms, although further studies are required. While our analysis focused on a specific architecture, we expect the results to be more broadly applicable. Other architectures using programmable nonlinearities may also see even more pronounced improvements.
We outline several open questions for future investigation. An important question involves the constituent optical blocks used to construct programmable quantum optical circuits. Specifically, it is of interest to explore how the choice of linear and nonlinear optical components affects the performance of QONNs. The NMZI block used in this work consists of two-mode linear optical DCs, but a variety of multimode couplers could be used to design other types of building blocks [43]. Additionally, while we focused on Kerr-like nonlinearity, there are many other nonlinearities that may act in more complex ways, including various multimode nonlinearities. These potentially accessible nonlinearities could facilitate the construction of more expressive quantum optical circuits by providing more sophisticated multivariate elements.
Another question pertains to methods for designing variational ansatze tailored to specific classes of problems, similar to gate-based variational circuits. In gate-based implementations of variational quantum algorithms, effectively solving a problem often requires designing a specific circuit architecture that accounts for problem specifics such as symmetries or conservation laws (references needed). This could also apply to the quantum optical platform, although methods from gate-based models cannot be directly applied. Therefore, research in this direction is necessary.
@@ Appendix A Transformation of the NMZI block
We consider a two-mode block shown in Fig. 1c which we call nonlinear MZI (NMZI). The NMZIs are constructed from the balanced directional couplers (DCs) characterized by the transfer matrices $U_{\text{DC}}=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}1&1\\ 1&-1\end{array}\right)$, and two programmable nonlinear sign-shift gates (NS-gates) placed in the arms between the DCs. The equation for NMZI contained two nonlinear elements  (1) with parameters $\chi_{1}$ and $\chi_{2}$ and linear phase shifter $\varphi$ for any input Fock state $|\psi_{in}\rangle=|n,m\rangle$ can be written as
$\begin{split}|\psi_{out}\rangle=\hat{U}_{\text{NMZI}}|n,m\rangle=\left(\dfrac{1}{2}\right)^{n+m}\sum_{p=0}^{n+m}\sum_{j=0}^{p}\begin{pmatrix}n\\ j\end{pmatrix}\times\\ \begin{pmatrix}m\\ p-j\end{pmatrix}(-1)^{m-p+j}e^{ip(p-1)\frac{\chi_{1}}{2}}e^{iq(q-1)\frac{\chi_{2}}{2}}e^{i\varphi p}\times\\ \sum_{\tilde{p}=0}^{n+m}\sum_{k=0}^{\tilde{p}}\begin{pmatrix}p\\ k\end{pmatrix}\begin{pmatrix}q\\ \tilde{p}-k\end{pmatrix}(-1)^{q-\tilde{p}+k}\dfrac{\sqrt{\tilde{p}!\tilde{q}!}}{\sqrt{n!m!}}|\tilde{p},\tilde{q}\rangle,\end{split}$ (11)

### v2

@@ II.2 Bell-state discrimination
$|\Phi^{(\pm)}\rangle=\frac{|0\rangle_{L1}|0\rangle_{L2}+|1\rangle_{L1}|1\rangle_{L2}}{\sqrt{2}}=\frac{|1010\rangle\pm|0101\rangle}{\sqrt{2}},$
$|\Psi^{(\pm)}\rangle=\frac{|0\rangle_{L1}|1\rangle_{L2}+|1\rangle_{L1}|0\rangle_{L2}}{\sqrt{2}}=\frac{|1001\rangle\pm|0110\rangle}{\sqrt{2}},$ (8)
where the right-hand side are the Bell-states encoded by $N=2$ photons and $M=4$ modes. The discriminator can transform each of the Bell-states $|\Psi^{(in)}_{j}\rangle\in\{|\Phi^{(+)}\rangle,|\Phi^{(-)}\rangle,|\Psi^{(+)}\rangle,|\Psi^{(-)}\rangle\}$ into a multimode Fock state $|\Psi^{(out)}_{j}\rangle\in\{|1010\rangle,|1001\rangle,|0110\rangle,|0101\rangle\}$. In turn, measuring these Fock states by photodetection unambiguously tells about a corresponding input state.
@@
$|\Theta^{(\pm)}\rangle=\frac{|0011\rangle\pm|1100\rangle}{\sqrt{2}},$ (9)
which are equally useful. Although states (9) do not respect the dual-rail encoding, they can still serve as a maximally entangled resource that fueled a quantum algorithm, provided cheap linear transformation is performed. Namely, to convert these states into any of the Bell states, one can apply simple two-mode operations, implemented deterministically by linear optics [43]. Therefore, in the following we consider the states (9) on par with the Bell states (II.2). We demand that the discriminator convert the input states $|\Theta^{(+)}\rangle$ and $|\Theta^{(-)}\rangle$ into $|\Psi^{(out)}_{5}\rangle=|0011\rangle$, $|\Psi^{(out)}_{6}\rangle=|1100\rangle$, respectively.
@@
Figure 8: Bell-state discrimination performed by the QONNs: state infidelity $1-F_{disc}$ as a function of circuit depth $D_{\text{NL}}$ for four- (a) and six-state discrimination (b) at different static phase-shift biases $\varphi_{b}=0$ of the QONN programmed by nonlinearities (solid curves). The dependencies corresponding to the QONN programmed by linear optics is plotted by the dashed curves. Each point is the best result of the optimization algorithm run $50$ times.
Figure 9: Distribution of the parameters of the trained QONNs: values of the NMZI nonlinearities (designated by the nonlinear layer index) and the MZI phase-shifts (designated by ”mzi” label) of the four-state (a) and six-state (b) discriminator at $\varphi_{b}=0$ with single-qubit corrections. The parameters correspond to the QONNs with $D_{\text{NL}}=1$ (a) and $D_{\text{NL}}=5$ (b) (see Fig. 8).
@@
The results suggest that the NMZI meshes, both with and without single-qubit corrections, can perform state discrimination provided proper phase-shift biases are set. The lowest depths at which the QONNs implement perfect discriminators are achieved at $\varphi_{b}=0$, corresponding to the identity operation of the NMZI in the linear regime. Specifically, depths of $D_{\text{NL}}=1$ and $D_{\text{NL}}=5$ are sufficient for the four- and six-state discrimination, respectively (with single-qubit corrections). A phase-shift of $\varphi_{b}=\pi$ (corresponding to the mode-swapping operation in the linear regime of NMZI) also works but requires much larger depths: $D_{\text{NL}}=3$ and $D_{\text{NL}}=11$ for the four- and six-state discrimination, respectively.
Training only the variational core without the input and output MZIs works as well, however, this comes at increasing the depths even further. In Fig. 8 we plotted the infidelities without the single-qubit corrections for $\varphi_{b}=0$ and $\varphi_{b}=\pi$. For these biases, it is possible to achieve convergence of the optimization algorithm only for $\varphi_{b}=\pi$, because NMZI mode swapping operation allows to redistribute the photons between modes. The minimal circuit depths in this case are $D_{\text{NL}}=19$ and $D_{\text{NL}}=23$.
@@
We summarize the relevant parameters of the QONN performing the state discrimination task in Table II.2. Notably, our QONN with adjustable nonlinearities is less demanding than QONNs with static nonlinearities, as evident from Fig. 8. Specifically, the QONN trained with linear optics requires $D_{\text{LO}}=10$ and $D_{\text{LO}}=20$ programmable linear phase-shift layers to achieve high-quality discrimination of the four- and six-state systems, respectively. Our QONN can do the task using only $D_{\text{NL}}=1$ and $D_{\text{NL}}=5$ programmable nonlinear phase-shift layers (at $\varphi_{b}=0$). Correspondingly, our QONN requires fewer adjustable nonlinearities than its counterpart trained with linear optics to find the optimal circuits.
Table 3: Nonlinear $D_{\text{NL}}$ and linear circuit depth $D_{\text{LO}}$ and parameter count $P$ of the QONNs peforming Bell-state discrimination (from the results shown in Fig. 8).
@@
cell11 = r=2, vlines, $D_{\text{NL}}$ $D_{\text{LO}}$ $P$
$4$-/$6$-state $4$-/$6$-state $4$-/$6$-state
NL ($\varphi_{b}=0$) 1/5 0 2/14
NL ($\varphi_{b}=\pi$) 3/11 0 8/32
NL ($\varphi_{b}=\pi$,no MZIs) 19/23 0 56/68
LO 1/3 10/20 24/48
@@ II.3 Variational quantum algorithms
We next consider QONN performing a variational quantum algorithm (VQA). VQAs enable solving certain types of computational problems by mapping them into corresponding Hamiltonians $\hat{H}$, whose ground state encodes the solutions. In order to find a ground state by VQA, one prepares parametrized trial states $|\Psi(\bm{\chi},\bm{\theta})\rangle$ on a quantum computer and estimates the Hamiltonian expectation value by measuring $E(\bm{\chi},\bm{\theta})=\langle\psi^{(dr)}(\bm{\chi},\bm{\theta})|\hat{H}|\psi^{(dr)}(\bm{\chi},\bm{\theta})\rangle$ — the cost function to be optimized by the classical computer, the global minimum $E_{\text{min}}\leq{}E(\bm{\chi},\bm{\theta})$ corresponds to a solution. Here, $|\psi^{(dr)}(\bm{\chi},\bm{\theta})\rangle$ designates the dual rail part of the output state $|\Psi(\bm{\chi},\bm{\theta})\rangle$. In contrast to implementing specific quantum operations, such as those described above, solving problems with the variational quantum eigensolver (VQE) does not require the exact generation of a specified state. VQE may provide good approximate solutions for a class of prepared quantum states that are not necessarily coinciding with the true ground state. Despite this relaxed complexity, solving practical problems by VQA is coupled with preparation of large-qubit quantum states, which, on the one hand, is too sophisticated due to the limitations of the current NISQ devices and, on the other hand, hard to optimize because of the unsolved issue of barrens plateau [38]. Therefore, developing less demanding quantum architectures for VQE is required.
@@
where $\hat{X}_{i},\hat{Y}_{i},\hat{Z}_{i}$ are the Pauli operators describing spin with index $i$. A particular Hamiltonian is specified by the interactions between the spins described by interaction coefficients $J_{ij}^{(\mu)}$ ($\mu=x,y,z$) with interaction graphs $G^{(\mu)}$ and external fields $h_{i}$. Given specified interactions and external fields, VQA boils down to finding the Hamiltonian ground state energy using an optimization algorithm run on a classical computer. Depending on the interaction graphs of the Hamiltonians, various quantum systems with different properties can be catched [48, 49]
@@
Figure 11: VQE of the Heisenberg Hamiltonian performed by the QONNs: energy error $\Delta{}E=E_{\text{exact}}-E_{\text{qonn}}$ as a function of depth $D_{\text{NL}}$ of the QONN programmed by nonlinearities at different values the NMZI phase-shift bias $\varphi_{b}$ (solid curves) and the dependence for the QONN programmed by linear optics (dashed curve). Each dependence correspond to $15$ Hamiltonians, obtained by randomly generating $J_{1}$, $J_{2}$ and $h_{i}$ from a uniform distribution in the range from 00 to $1$. For each value of $D_{\text{NL}}$ and each random Hamiltonian, the best $\Delta{}E$ from $50$ optimization runs was taken into the distribution.
Specifically, we sampled a number of Hamiltonians by generating values of $J_{1}$, $J_{2}$ and $h_{i}$ ($i=\overline{1,5}$) randomly from the uniform distribution over in the range $[0,1]$. Here, we quantify the infidelity by the energy error $\Delta{}E=E_{\text{exact}}-E_{\text{qonn}}$, where $E_{\text{qonn}}$ is the energy obtained by training/optimizing a QONN, $E_{\text{exact}}$ is its exact value obtained by the diagonalization of the Hamiltonian matrix using a standard library (Python SciPy). The obtained dependencies of error on the circuit depth $D_{\text{NL}}$ is shown in Fig. 11. As before, not all static phase-shift values $\varphi_{b}$ enabled find as a high-accuracy solution. As we see, the value $\varphi_{b}=0$ (corresponding to the mode swapping operation in the linear regime of NMZI) does not yield satisfactory quality. At the same time, all the rest of the values work almost equally good.
According to Fig. 11, it requires minimum $D_{\text{NL}}=7$ programmable nonlinear layers for the best case of $\varphi_{b}=\pi/2$ for our QONN to be able to find high-quality solutions, while the QONN programmed by linear optics requires $D_{\text{LO}}=22$ programmable linear layers. This suggests that our QONN is advantageous.
Table 4: Nonlinear $D_{\text{NL}}$ and linear circuit depth $D_{\text{LO}}$ and parameter count $P$ in the QONNs peforming VQE of the Heisenberg Hamiltonian (see Fig. 11). NL and LO are the QONNs with adjustable nonlinearities and the counterparts trained by linear optics, respectively.
@@
vlines, $D_{\text{NL}}$ $D_{\text{LO}}$ $P$
NL ($\varphi_{b}=\pi/2$) 7 0 62
LO 1 22 180
It is worth noting that the circuit depth, and therefore the number of adjustable parameters in QONNs, can be further reduced by using different building blocks. Specifically, the ability of a QONN layer to implement non-local interactions may be more advantageous for certain problems compared to the two-mode interactions of NMZI layers. Our proposal is not confined to this specific QONN architecture. We suggest exploring the use of programmable nonlinearities in other architectures that may be developed in the future.
@@ II.4 Training complexity
Figure 12: Training the QONNs for preparation of the $4$-qubit GHZ-state: a) the set of infidelities to which the optimization algorithm has converged in $30$ optimization runs and the corresponding number of evaluations of the cost function; b) examples of convergence curves in $10$ optimization runs. The parameters of the QONNs correspond to Fig. 5e) and Fig. 5f) at $D_{\text{NL}}=9$ and $D_{\text{NL}}=3$, respectively.
@@ III Discussion and Outlook
We have proposed a novel approach for constructing parametrized quantum optical circuits using programmable nonlinearities. Using NMZI meshes as an example of such programmable circuits, we have demonstrated that high-quality operation can be achieved by training these nonlinearities. The key insight of our work is that relying solely on programmable linear optical interference is suboptimal for implementing QONNs. Our approach, which incorporates programmable nonlinearities, significantly reduces the number of layers of programmable elements, thereby decreasing errors from circuit components and resulting in more compact designs. Additionally, variational circuits can benefit from a lower number of adjustable parameters, allowing for more computationally efficient optimization algorithms, although further studies are required. Although our analysis focused on a specific architecture, we expect the results to be more broadly applicable. Other architectures using programmable nonlinearities may also see even more pronounced improvements. Also, extending our approach to the continuous variable quantum neural networks [52] may yield similar improvements.
We outline several open questions for future investigation. An important question involves the constituent optical blocks used to construct programmable quantum optical circuits. Specifically, it is of interest to explore how the choice of linear and nonlinear optical components affects the performance of QONNs. The NMZI block used in this work consists of two-mode linear optical DCs, but a variety of multimode couplers could be used to design other types of building blocks. Additionally, while we focused on Kerr-like nonlinearity, there are many other nonlinearities that may act in more complex ways, including various multimode nonlinearities. These potentially accessible nonlinearities together with multimode linear optical elements could facilitate the construction of more expressive and/or more resource-efficient quantum optical circuits by providing more sophisticated multivariate elements.
Another question pertains to methods for designing variational ansatze tailored to specific classes of problems, similar to gate-based variational circuits. In gate-based implementations of variational quantum algorithms, effectively solving a problem often requires designing a specific circuit architecture that accounts for problem specifics such as symmetries or conservation laws. This could also apply to the quantum optical platform, although methods from gate-based models cannot be directly used. Therefore, research in this direction is necessary.
@@ Appendix A Transformation of the NMZI block
We consider a two-mode block, shown in Fig. 1c which we call nonlinear MZI (NMZI). The NMZIs are constructed from the balanced directional couplers (DCs) characterized by the transfer matrices $U_{\text{DC}}=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}1&1\\ 1&-1\end{array}\right)$, and two programmable nonlinear sign-shift gates (NS-gates) placed in the arms between the DCs. The equation for NMZI contained two nonlinear elements $\hat{\text{NS}}_{m}(\chi)=\exp\left(i\frac{\chi}{2}(\hat{a}_{m}^{\dagger})^{2}\hat{a}_{m}^{2}\right)$ (1) with parameters $\chi_{1}$ and $\chi_{2}$ and linear phase shifter $\varphi$ for any input Fock state $|\psi_{in}\rangle=|n,m\rangle$ can be written as
$\begin{split}|\psi_{out}\rangle=\hat{U}_{\text{NMZI}}|n,m\rangle=\left(\dfrac{1}{2}\right)^{n+m}\sum_{p=0}^{n+m}\sum_{j=0}^{p}\begin{pmatrix}n\\ j\end{pmatrix}\times\\ \begin{pmatrix}m\\ p-j\end{pmatrix}(-1)^{m-p+j}\times\\ \times e^{ip(p-1)\frac{\chi_{1}}{2}}e^{iq(q-1)\frac{\chi_{2}}{2}}e^{i\varphi p}\times\\ \sum_{\tilde{p}=0}^{n+m}\sum_{k=0}^{\tilde{p}}\begin{pmatrix}p\\ k\end{pmatrix}\begin{pmatrix}q\\ \tilde{p}-k\end{pmatrix}(-1)^{q-\tilde{p}+k}\dfrac{\sqrt{\tilde{p}!\tilde{q}!}}{\sqrt{n!m!}}|\tilde{p},\tilde{q}\rangle,\end{split}$ (12)